# Optimal Nuisance Function Tuning for Estimating a Doubly Robust Functional under Proportional Asymptotics

### v1

@@ 3.1 Bias correction
3.1.1 Integral based estimator
@@
where $F_{\rm MP}$ is the famous Marchenko-Pasteur distribution (Section 3 of [24]), which arises as the limiting spectral distribution (LSD) of the eigenvalues of the sample covariance matrix ($\hat{\Sigma}_{1}$ or $\hat{\Sigma}_{2}$). The above sketch indicates that $\mathbb{E}[\hat{\alpha}(\lambda_{1})^{\top}\hat{\beta}(\lambda_{2})]\approx\textstyle(\alpha_{0}^{\top}\beta_{0})g^{\rm INT}_{\rm 3sp}(\lambda_{1},\lambda_{2})$, i.e. the ridge-regression based plug-in estimates approximate $(\alpha_{0}^{\top}\beta_{0})$ up to a scalar function $g^{\rm INT}_{\rm 3sp}(\lambda_{1},\lambda_{2})$. This scalar function, however, is deterministic, as it does not depend on any unknown parameters. Therefore, it naturally motivates a simple debiasing procedure by inverting the asymptotic bias:
@@ 3.2 Limiting variance
It is worth noting that the optimal choice of $(\lambda_{1},\lambda_{2})$ (which is the minimizer of Equation (3.10)) differs significantly from the prediction optimal choice of $(\lambda_{1},\lambda_{2})$, as illustrated by our simulation in Section 4. This departs from the traditional DML framework, where the nuisance parameters are typically estimated using prediction-optimal methods. However, in the proportional asymptotic regime, such prediction-optimal choices do not necessarily yield the best performance for inference. As a result, it is more appropriate to directly minimize the asymptotic variance rather than predict optimal choices to achieve a more accurate inference/narrower confidence interval.
@@ 4 Simulations
We generated $10,000$ independent data sets, each consisting of iid copies of $(A,X,Y)$, and assessed the performance of estimators of the ECC through simulations under two sample splitting regimes: two-split with a total sample size $N=1,000$, and three-split with $N=1,500$, with splits of size 500. Specifically, we consider a random vector $X$ of dimension $p$, generated as $X\sim\mathcal{N}(0,I_{p})$. Then, we generated $A$ and $Y$ with $\rho=0.5$ by Equation (2.1).
For each sample splitting regime, we assess the performance of estimators of the ECC with dimensionalities of $p=250$ (i.e., $c=p/n=0.25$) and $p=1000$ (i.e., $c=p/n=2$). The numerical estimation of bias correction constants was conducted using Monte Carlo with $10,000$ iterations. Each estimator was applied across 100 distinct values of $\lambda$, spanning a grid from 0.05 to 10.
Our analysis primarily focused on the influence of the tuning parameter $\lambda$ on the variance of the debiased estimators, including the Integral-based estimator, the Newey-Robins estimator, and the Doubly robust estimator. Results for $p=1000$ (i.e., $c=2$) are presented in Figure 1, while technical details for the $p=250$ setting (i.e., $c=0.5$) and bias comparisons between original and debiased estimators are provided in the Appendix. Interestingly, we demonstrate that prediction-optimal choices of the tuning parameters fail to yield statistically optimal estimates of ECC, and one should rather aim at optimizing the variance of the ECC directly and carefully to account for the type of sample splitting and estimator chosen. We note that in the two-split scenario, the variance blows up at values around $\lambda=1.48$ due to the bias-correction constant. However, this does not affect our primary conclusions, and this phenomenon will be further investigated.
Figure 1: Variances of debiased estimators as functions of tuning parameter $\lambda$. Top row: Two-split (N = 1000 total, split into two subsamples of 500 each). Bottom row: Three-split (N = 1500 total, split into three subsamples of 500 each). Estimator (left to right): integral-based plug-in estimator, the Newey-Robins plug-in estimator, and the doubly robust estimator. Red line indicates the optimal $\lambda$ for estimator variance; Blue line is for prediction-optimal $\lambda$. The figures shown here are zoomed-in views; complete versions are included in the Appendix.
@@ 5 Conclusion
In this paper, we have explored the intricate interplay between sample splitting, estimator choice, and optimal tuning of nuisance functions for estimating a popular doubly robust functional of interest. The philosophy we espouse in this regard can be best understood against the DML framework, where nuisance functions are typically estimated for prediction optimality for downstream efficient estimation of a low-dimensional functional that depends on the nuisance functions. Specifically, by considering a popular functional of interest, we explore the subtleties of this discourse when no consistent nuisance functions exist. While considering ECC, one such challenging regime arises when the dimension of the problem is proportional to the sample size and our results uncover a precise asymptotic picture of the interplay between sample splitting, choice of estimator, and optimal tuning of ridge regularized nuisance function in this regard under a stylized linear model for the nuisance functions. Our results explore this regime under linear models for nuisance functions and homoscedastic correlation. A limitation of our analysis pertains to the assumption of the gaussianity of either the errors or covariates for variance calculation – an aspect that remains to be explored through the lens of universality principles. It is, however, worth noting that, unlike the operational domain of the DML proposal, we consider a regime where asymptotic optimality for ECC has not yet been established, and the literature has been active in constructing estimators that navigate specific contours of the problem difficulty, such as debiasing regularized estimators of the nuisance functions. Indeed, the estimators studied here are not the only ones that achieve $\sqrt{n}$-consistency in this regime. Indeed, one can completely bypass nuisance function estimation through a method of moments type approach [26]. One can also envision appealing to more sophisticated debiasing techniques when generic convex regularization, such as Lasso, is employed to estimate the nuisance functions [25] and derive results analogous to those presented here. We believe that the template of exploration presented here might serve as a framework for such future endeavors. Specifically, the availability of a large class of sample split sensitive optimally tuned estimators of double robust functionals might shed light on eventual efficiency bounds for estimating the functional itself. A case study of ECC presented here hopefully serves as a compass for future conjectures in this journey.
@@
A.1 Proof for three-split
@@ Appendix F Prediction-Optimal Tuning Parameters
Appendix G Additional details and results for the simulation study
Here, we provide additional details and results of the simulation study. The additional results we present here serve the following purposes:
@@ Appendix G Additional details and results for the simulation study
    We illustrate the asymptotic bias of the integral-based estimator, Newey Robins estimator, and doubly robust estimator and we verify that the debiased versions of these estimators are asymptotically unbiased.

### v2

@@ 3.1 Bias Correction
3.1.1 Integral-Based Estimator
@@
where $F_{\rm MP}$ is the famous Marchenko-Pastur distribution (Section 3 of [24]), which arises as the limiting spectral distribution (LSD) of the eigenvalues of the sample covariance matrix ($\hat{\Sigma}_{1}$ or $\hat{\Sigma}_{2}$). The above sketch indicates that $\mathbb{E}[\hat{\alpha}(\lambda_{1})^{\top}\hat{\beta}(\lambda_{2})]\approx\textstyle(\alpha_{0}^{\top}\beta_{0})g^{\rm INT}_{\rm 3sp}(\lambda_{1},\lambda_{2})$, i.e. the ridge-regression based plug-in estimates approximate $(\alpha_{0}^{\top}\beta_{0})$ up to a scalar function $g^{\rm INT}_{\rm 3sp}(\lambda_{1},\lambda_{2})$. This scalar function, however, is deterministic, as it does not depend on any unknown parameters. Therefore, it naturally motivates a simple debiasing procedure by inverting the asymptotic bias:
@@ 3.2 Limiting Variance
It is worth noting that the optimal choice of $(\lambda_{1},\lambda_{2})$ (which is the minimizer of Equation (3.10)) differs significantly from the prediction-optimal choice of $(\lambda_{1},\lambda_{2})$ (derived in the Appendix), as illustrated by our simulation in Section 4. This departs from the traditional DML framework, where the nuisance parameters are typically estimated using prediction-optimal methods. However, in the proportional asymptotic regime, such prediction-optimal choices do not necessarily yield the best performance for inference. As a result, it is more appropriate to directly minimize the asymptotic variance rather than predict optimal choices to achieve a more accurate inference/narrower confidence interval.
In practice, the asymptotic variances can be estimated by using resampling techniques. We outline a parametric bootstrap approach to estimate the asymptotic variances in the Appendix, which we evaluate in our simulations.
@@ 4 Simulations
We assessed the performance of the ECC estimators through simulations under the different sample splitting strategies. We considered the size of the splits to be $500$, making the total sample size $N=1000$ for the two sample splitting case and $N=1500$ for the three sample splitting case. For each sample splitting case, we set $p=250$ (i.e., $c=0.25)$ and $p=1000$ (i.e., $c=2)$. In each case, we generated $10,000$ independent data sets, each consisting of iid copies of $(A,X,Y)$. Specifically, we consider a random vector $X$ of dimension $p$, generated as $X\sim\mathcal{N}(0,I_{p})$. Then, we generated $A$ and $Y$ with $\rho=0.5$ by Equation (2.1). We applied the integral-based estimator, the Newey-Robins estimator, and the doubly robust estimator, where we set $\lambda_{1}=\lambda_{2}=\lambda$ and set the bias correction constants by using Monte Carlo with $10,000$ iterations. Each estimator was applied across 100 distinct values of $\lambda$, spanning a grid from 0.05 to 10.
We investigated the impact of the tuning parameter $\lambda$ on the asymptotic variance of the debiased estimators. Results for $p=1000$ (i.e., $c=2$) are presented in Figure 1, while results for the $p=250$ setting (i.e., $c=0.5$) are provided in the Appendix. Interestingly, we demonstrate that prediction-optimal choices of the tuning parameters fail to yield statistically optimal estimates of ECC, and one should rather aim at optimizing the variance of the ECC directly and carefully to account for the type of sample splitting and estimator chosen. We note that in the two-split scenario, the variance blows up at values around $\lambda=1.48$ due to the bias-correction constant diverging. However, this does not affect our primary conclusions, as our goal is to select the value of $\lambda$ that minimizes the asymptotic variance.
Due to space constraints, we present additional details of the simulations and our other analyses in the Appendix. In brief, we found that the debiased estimators had negligible bias (unlike the original estimators) and that the parametric bootstrap approach often performed well for estimating the asymptotic variance.
Figure 1: Variances of debiased estimators as functions of tuning parameter $\lambda$. Top row: Two-split ($N=1000$ total, split into two subsamples of 500 each). Bottom row: Three-split ($N=1500$ total, split into three subsamples of 500 each). Red line indicates the optimal $\lambda$ for estimator variance; Blue line is for prediction-optimal $\lambda$. The figures shown here are zoomed-in views; complete versions are included in the Appendix.
@@ 5 Conclusion
In this paper, we have explored the intricate interplay among sample splitting, estimator choice, and optimal tuning of nuisance functions for estimating the expected conditional covariance (ECC), a popular doubly robust functional in the literature on causal inference and missing data. The philosophy we espouse in this regard can be best understood within the DML framework, where typically prediction-optimal nuisance functions are used for downstream efficient estimation of a low-dimensional functional that depends on them. In this paper, we explore the subtleties of this discourse when no consistent nuisance functions exist. In particular, for ECC, this challenging regime arises when the dimension of the problem is proportional to the sample size, and our results uncover a precise asymptotic picture of the interplay between sample splitting, choice of estimator, and optimal tuning of the regularized nuisance function in this regard under a stylized linear model for the nuisance functions. Our analysis reveals an interesting phenomenon: prediction-optimal tuning parameters are not necessarily the best choices for estimating ECC, in terms of its asymptotic variance (and consequently the length of the confidence interval), which differs from standard DML wisdom.
While this is the first work to analyze ECC in the proportional over-parameterized regime, our framework invites several natural extensions. For example, our analysis assumes the Gaussianity of either the errors or the covariates for variance calculation – an assumption that remains to be explored through the lens of universality principles. It is, however, worth noting that, unlike the operational domain of the DML proposal, we consider a regime where asymptotic optimality for ECC has not yet been established, and the literature has been active in constructing estimators that navigate specific contours of the problem difficulty, such as debiasing regularized estimators of the nuisance functions. Indeed, the estimators studied here are not the only ones that achieve $\sqrt{n}$-consistency in this regime; one can completely bypass nuisance function estimation using a method-of-moments approach [26]. One can also envision appealing to more sophisticated debiasing techniques when generic convex regularization, such as Lasso, is employed to estimate the nuisance functions [25] and derive results analogous to those presented here. In this regime, where an efficiency theory is unavailable, the choice of estimator that minimizes asymptotic variance is a priori unclear. We believe that the template of exploration presented here might serve as a framework for such future endeavors. Specifically, the availability of a large class of sample-split-sensitive, optimally tuned estimators of double-robust functionals might shed light on eventual efficiency bounds for estimating the functional itself. A case study of ECC presented here hopefully serves as a compass for future conjectures in this journey.
Acknowledgments and Disclosure of Funding
Rajarshi Mukherjee was supported by NSF CAREER award 8529216-01. Debarghya Mukherjee was supported by NSF DMS-2515787. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NSF.
@@
A.1 Proof of Three-Split
@@ Appendix F Prediction-Optimal Tuning Parameters
Appendix G Parametric Bootstrap
We use a parametric bootstrap approach to estimate the large-sample variance of our debiased estimators. For illustration purpose, we consider $\hat{\theta}_{\rm 3sp}^{\rm db,INT}$, which is defined as:
$\hat{\rho}=\hat{\theta}_{\rm 3sp}^{\rm db,INT}=\frac{1}{n}\sum_{i\in\mathcal{D}_{3}}Y_{i}A_{i}-\ \frac{\hat{\alpha}(\lambda_{1})^{\top}\hat{\beta}(\lambda_{2})}{g^{\rm INT}_{3sp}(\lambda_{1},\lambda_{2})}$
where $g^{\rm INT}_{\rm 3sp}(\lambda_{1},\lambda_{2})=\Big(\int\frac{x\,dF_{\rm MP}(x)}{x+\lambda_{1}}\Big)\Big(\int\frac{x\,dF_{\rm MP}(x)}{x+\lambda_{2}}\Big)$ and $F_{\rm MP}$ denotes the Marchenko–Pastur law with aspect ratio $c=p/n$.
Bootstrap algorithm (3-split).
Given $\hat{\rho}$ and the ridge estimators $(\hat{\alpha}(\lambda_{1}),\hat{\beta}(\lambda_{2}))$, we define the following:
@@ Appendix G Parametric Bootstrap
    Define the constants $(c_{1},d_{1},c_{2},d_{2},c_{3})$ as:
    $c_{j}$ $=\int\frac{x^{2}}{(x+\lambda_{j})2}\ dF_{\rm MP}(x),\qquad j\in\{1,2\}$
    $d_{j}$ $=(p/n)\int\frac{x}{(x+\lambda_{j})^{2}}\ dF_{\rm MP}(x),\qquad j\in\{1,2\}$
    $c_{3}$ $=\int\frac{x}{x+\lambda_{1}}\ dF_{\rm MP}(x)\ \int\frac{x}{x+\lambda_{2}}\ dF_{\rm MP}(x)\,.$
    Then define two transformed regression coefficients $\tilde{\alpha}$ and $\tilde{\beta}$ as:
    $\tilde{\alpha}$ $=\frac{\hat{\alpha}(\lambda_{1})}{\|\hat{\alpha}(\lambda_{1})\|_{2}}\sqrt{\frac{\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}-d_{1}}{c_{1}}}$
    $\tilde{\beta}$ $=\frac{\hat{\alpha}(\lambda_{1})}{\|\hat{\alpha}(\lambda_{1})\|_{2}}\left(\frac{(\hat{\alpha}(\lambda_{1})^{\top}\hat{\beta}(\lambda_{2}))}{c_{3}}\frac{1}{\sqrt{\frac{\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}-d_{1}}{c_{1}}}}\right)\ +\mathbf{z}\sqrt{\frac{\|\hat{\beta}(\lambda_{2})\|_{2}^{2}-d_{2}}{c_{2}}-\left(\frac{(\hat{\alpha}(\lambda_{1})^{\top}\hat{\beta}(\lambda_{2}))}{c_{3}}\frac{1}{\sqrt{\frac{\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}-d_{1}}{c_{1}}}}\right)^{2}}$
    where $\mathbf{z}$ is some unit vector perpendicular to $\hat{\alpha}(\lambda_{1})$.
@@
    Generate $X^{(b)}_{1},\dots,X^{(b)}_{n}\sim\mathcal{N}(0,I_{p})$.
  3. 3.
    Generate $({\epsilon}^{(b)}_{1},\eta^{(b)}_{1}),\dots,({\epsilon}^{(b)}_{n},\eta^{(b)}_{n})\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\hat{\rho}\\ \hat{\rho}&1\end{pmatrix}\right)$.
  4. 4.
    Set $A^{(b)}_{i}=X^{(b)^{\top}}_{i}\tilde{\alpha}+{\epsilon}^{(b)}_{i}$ and $Y^{(b)}_{i}=X^{(b)^{\top}}_{i}\tilde{\beta}+\eta^{(b)}_{i}$ for $1\leq i\leq n$.
  5. 5.
    Estimate $\hat{\theta}^{(b),\rm db,INT}_{3sp}$ using $\{(X_{i}^{(b)},Y_{i}^{(b)},A_{i}^{(b)})\}_{1\leq i\leq n}$.
  6. 6.
    Repeat Step 2-5 multiple (say $B$) times and return:
    $\widehat{{\sf var}}=\frac{1}{B}\sum_{b=1}^{B}\left(\hat{\theta}^{(b),\rm db,INT}_{3sp}\right)^{2}-\left(\frac{1}{B}\sum_{b=1}^{B}\hat{\theta}^{(b),\rm db,INT}_{3sp}\right)^{2}\,.$
[…]
$\lim_{n\uparrow\infty}\ n\times{\sf var}\left(\hat{\theta}^{(b),\rm db,INT}_{3sp}\right)=\sigma^{2}(\rho,\varrho,u,v,c,\lambda_{1},\lambda_{2})\,.$
where $u=\lim_{p\uparrow\infty}\|\alpha\|^{2}_{2},\ v=\lim_{n\uparrow\infty}\|\beta\|^{2}_{2},\ \varrho=\lim_{n\uparrow\infty}\alpha^{\top}\beta$ and $c$ is the limiting value of $(p/n)$. On the other hand, if we increase the bootstrap sample size $B$, then $\hat{\sf var}$ converges to the conditional variance of $\hat{\theta}_{\rm 3sp}^{\rm db,INT}$ given the original data, i.e.,
$\lim_{B\uparrow\infty}\widehat{{\sf var}}={\sf var}\left(\hat{\theta}^{(b),\rm db,INT}_{3sp}\mid\mathcal{D}\right)\,.$
Therefore, the bootstrap variance estimation will be consistent if we can show that:
$\lim_{n\uparrow\infty}{\sf var}\left(\hat{\theta}^{(b),\rm db,INT}_{3sp}\mid\mathcal{D}\right)\overset{\text{a.s.}}{=}\sigma^{2}(\rho,\varrho,u,v,c,\lambda_{1},\lambda_{2})\,.$ (G.1)
However there is a small issue here: we know that neither $\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}$ nor $\hat{\|}\beta(\lambda_{2})\|_{2}^{2}$ gives us a consistent estimate of $u$ of $v$. In particular we can show that:
$\lim_{n\uparrow\infty}\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}$ $\overset{a.s.}{=}c_{1}u+d_{1}$
$\lim_{n\uparrow\infty}\|\hat{\beta}(\lambda_{2})\|_{2}^{2}$ $\overset{a.s.}{=}c_{2}u+d_{2}$
$\lim_{n\uparrow\infty}\hat{\alpha}(\lambda_{1})^{\top}\hat{\beta}(\lambda_{2})$ $\overset{a.s.}{=}c_{3}\varrho\,.$
where $(c_{1},d_{1},c_{2},d_{2},c_{3})$ is same defined in the Step 1 of our bootstrap algorithm. As a consequence, one can show that:
$\lim_{n\uparrow\infty}{\sf var}\left(\hat{\theta}^{(b),\rm db,INT}_{3sp}\mid\mathcal{D}\right)\overset{\text{a.s.}}{=}\sigma^{2}(\rho,c_{3}\varrho,c_{1}u+d_{1},c_{2}v+d_{2},c,\lambda_{1},\lambda_{2})$
[…]
$\lim_{n\uparrow\infty}\|\tilde{\alpha}\|_{2}^{2}$ $=\lim_{n\uparrow\infty}\left\|\frac{\hat{\alpha}(\lambda_{1})}{\|\hat{\alpha}(\lambda_{1})\|_{2}}\sqrt{\frac{\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}-d_{1}}{c_{1}}}\right\|_{2}^{2}$
$=\lim_{n\uparrow\infty}\frac{\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}-d_{1}}{c_{1}}$
$=\frac{c_{1}u+d_{1}-d_{1}}{c_{1}}=u\,.$
$\lim_{n\uparrow\infty}\|\tilde{\beta}\|_{2}^{2}$ (G.2)
$=\lim_{n\uparrow\infty}\left\|\frac{\hat{\alpha}(\lambda_{1})}{\|\hat{\alpha}(\lambda_{1})\|_{2}}\left(\frac{(\hat{\alpha}(\lambda_{1})^{\top}\hat{\beta}(\lambda_{2}))}{c_{3}}\frac{1}{\sqrt{\frac{\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}-d_{1}}{c_{1}}}}\right)+\mathbf{z}\sqrt{\frac{\|\hat{\beta}(\lambda_{2})\|_{2}^{2}-d_{2}}{c_{2}}-\left(\frac{(\hat{\alpha}(\lambda_{1})^{\top}\hat{\beta}(\lambda_{2}))}{c_{3}}\frac{1}{\sqrt{\frac{\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}-d_{1}}{c_{1}}}}\right)^{2}}\right\|_{2}^{2}$ (G.3)
$=\lim_{n\uparrow\infty}\left[\left(\frac{(\hat{\alpha}(\lambda_{1})^{\top}\hat{\beta}(\lambda_{2}))}{c_{3}}\frac{1}{\sqrt{\frac{\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}-d_{1}}{c_{1}}}}\right)^{2}+\left\{\frac{\|\hat{\beta}(\lambda_{2})\|_{2}^{2}-d_{2}}{c_{2}}-\left(\frac{(\hat{\alpha}(\lambda_{1})^{\top}\hat{\beta}(\lambda_{2}))}{c_{3}}\frac{1}{\sqrt{\frac{\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}-d_{1}}{c_{1}}}}\right)^{2}\right\}\right]$ (G.4)
$=\lim_{n\uparrow\infty}\frac{\|\hat{\beta}(\lambda_{2})\|_{2}^{2}-d_{2}}{c_{2}}$ (G.5)
$=\frac{c_{2}v+d_{2}-d_{2}}{c_{2}}=v$ (G.6)
Last, but not least,
$\lim_{n\uparrow\infty}\tilde{\alpha}^{\top}\tilde{\beta}$
$=\lim_{n\uparrow\infty}\left\langle\frac{\hat{\alpha}(\lambda_{1})}{\|\hat{\alpha}(\lambda_{1})\|_{2}}\sqrt{\frac{\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}-d_{1}}{c_{1}}},\frac{\hat{\alpha}(\lambda_{1})}{\|\hat{\alpha}(\lambda_{1})\|_{2}}\left(\frac{(\hat{\alpha}(\lambda_{1})^{\top}\hat{\beta}(\lambda_{2}))}{c_{3}}\frac{1}{\sqrt{\frac{\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}-d_{1}}{c_{1}}}}\right)\right\rangle$
$=\lim_{n\uparrow\infty}\frac{(\hat{\alpha}(\lambda_{1})^{\top}\hat{\beta}(\lambda_{2}))}{c_{3}}=\varrho\,.$
[…]
For the other estimators in the three-split regime, we apply the same three-split bootstrap procedure described above. The only difference lies in the specific form of the estimators themselves (i.e., the analytical expressions for $\hat{\theta}^{\rm db,NR}_{\rm 3sp}$ and $\hat{\theta}^{\rm db,DR}_{\rm 3sp}$), while the underlying bootstrap mechanism and data-generating steps remain identical.
$\lim_{n\uparrow\infty}\ n\times{\sf var}\left(\hat{\theta}^{(b),\rm db,INT/NR/DR}_{3sp}\right)=\sigma^{2}(\rho,\varrho,u,v,c,\lambda_{1},\lambda_{2})\,.$
Since all three estimators share the same asymptotic structure and depend on the same limiting quantities $(u,v,\varrho)$. Thus, the consistency argument for the bootstrap variance estimation holds in exactly the same way for NR and DR as for the INT estimator.
G.1 Two-Split Parametric Bootstrap
With two splits, the norms obey the same limits, but the cross term changes. Let
$\hat{\alpha}(\lambda_{1})^{\top}\hat{\beta}(\lambda_{2})\ \xrightarrow{\rm a.s.}\ g_{1}^{\rm INT}(\lambda_{1},\lambda_{2})\,\varrho+g_{2}^{\rm INT}(\lambda_{1},\lambda_{2})\,\rho,$
where we have $g_{1,\rm 2sp}^{\rm INT}=\int\frac{x^{2}\,dF_{\rm MP}(x)}{(x+\lambda_{1})(x+\lambda_{2})},g_{2,\rm 2sp}^{\rm INT}(\lambda_{1},\lambda_{2})=\int\frac{cx\,dF_{\rm MP}(x)}{(x+\lambda_{1})(x+\lambda_{2})}$
$g_{1}^{\rm INT}(\lambda_{1},\lambda_{2})=\int\frac{x^{2}}{(x+\lambda_{1})(x+\lambda_{2})}\,dF_{\rm MP}(x),\qquad g_{2}^{\rm INT}(\lambda_{1},\lambda_{2})=\int\frac{c\,x}{(x+\lambda_{1})(x+\lambda_{2})}\,dF_{\rm MP}(x).$
Let $\hat{\mathbf{a}}:=\hat{\alpha}(\lambda_{1})/\|\hat{\alpha}(\lambda_{1})\|_{2}$ and let $\mathbf{z}$ be any unit vector orthogonal to $\hat{\mathbf{a}}$. Define
$\tilde{\alpha}$ $:=\hat{\mathbf{a}}\,\sqrt{\frac{\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}-d_{1}}{c_{1}}},$ (G.7)
$\tilde{\beta}$ $:=\hat{\mathbf{a}}\,\left(\frac{\hat{\alpha}(\lambda_{1})^{\top}\hat{\beta}(\lambda_{2})-g_{2}\hat{\rho}}{g_{1}}\;\frac{1}{\sqrt{\frac{\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}-d_{1}}{c_{1}}}}\right)\;+\;\mathbf{z}\,\sqrt{\frac{\|\hat{\beta}(\lambda_{2})\|_{2}^{2}-d_{2}}{c_{2}}-\left(\frac{\hat{\alpha}(\lambda_{1})^{\top}\hat{\beta}(\lambda_{2})-g_{2}\hat{\rho}}{g_{1}}\;\frac{1}{\sqrt{\frac{\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}-d_{1}}{c_{1}}}}\right)^{2}}.$
We rewrite
$g_{1}$ $=g_{1}^{\rm INT}(\lambda_{1},\lambda_{2})\quad g_{2}=g_{2}^{\rm INT}(\lambda_{1},\lambda_{2})$
$\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}\$$\xrightarrow{\text{a.s.}}\ c_{1}u+d_{1},\qquad\|\hat{\beta}(\lambda_{2})\|_{2}^{2}\ \xrightarrow{\text{a.s.}}\ c_{2}v+d_{2},\qquad\hat{\alpha}(\lambda_{1})^{\top}\hat{\beta}(\lambda_{2})\ \xrightarrow{\text{a.s.}}\ g_{1}\varrho+g_{2}\rho,$
$\lim_{n\uparrow\infty}\|\tilde{\alpha}\|_{2}^{2}$ $=\lim_{n\uparrow\infty}\left\|\frac{\hat{\alpha}(\lambda_{1})}{\|\hat{\alpha}(\lambda_{1})\|_{2}}\sqrt{\frac{\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}-d_{1}}{c_{1}}}\right\|_{2}^{2}$
$=\lim_{n\uparrow\infty}\frac{\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}-d_{1}}{c_{1}}=\frac{c_{1}u+d_{1}-d_{1}}{c_{1}}=u\,.$
$\lim_{n\uparrow\infty}\|\tilde{\beta}\|_{2}^{2}$ (G.8)
$=\lim_{n\uparrow\infty}\Bigg\|\frac{\hat{\alpha}(\lambda_{1})}{\|\hat{\alpha}(\lambda_{1})\|_{2}}\left(\frac{\hat{\alpha}(\lambda_{1})^{\top}\hat{\beta}(\lambda_{2})-g_{2}\,\hat{\rho}}{g_{1}}\,\frac{1}{\sqrt{\frac{\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}-d_{1}}{c_{1}}}}\right)+\mathbf{z}\sqrt{\frac{\|\hat{\beta}(\lambda_{2})\|_{2}^{2}-d_{2}}{c_{2}}-\left(\frac{\hat{\alpha}(\lambda_{1})^{\top}\hat{\beta}(\lambda_{2})-g_{2}\,\hat{\rho}}{g_{1}}\,\frac{1}{\sqrt{\frac{\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}-d_{1}}{c_{1}}}}\right)^{2}}\Bigg\|_{2}^{2}$ (G.9)
$=\lim_{n\uparrow\infty}\left[\left(\frac{\hat{\alpha}(\lambda_{1})^{\top}\hat{\beta}(\lambda_{2})-g_{2}\,\hat{\rho}}{g_{1}}\,\frac{1}{\sqrt{\frac{\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}-d_{1}}{c_{1}}}}\right)^{2}+\left\{\frac{\|\hat{\beta}(\lambda_{2})\|_{2}^{2}-d_{2}}{c_{2}}-\left(\frac{\hat{\alpha}(\lambda_{1})^{\top}\hat{\beta}(\lambda_{2})-g_{2}\,\hat{\rho}}{g_{1}}\,\frac{1}{\sqrt{\frac{\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}-d_{1}}{c_{1}}}}\right)^{2}\right\}\right]$ (G.10)
$=\lim_{n\uparrow\infty}\frac{\|\hat{\beta}(\lambda_{2})\|_{2}^{2}-d_{2}}{c_{2}}=\frac{c_{2}v+d_{2}-d_{2}}{c_{2}}=v\,.$ (G.11)
Last, but not least,
$\lim_{n\uparrow\infty}\tilde{\alpha}^{\top}\tilde{\beta}$ $=\lim_{n\uparrow\infty}\left\langle\frac{\hat{\alpha}(\lambda_{1})}{\|\hat{\alpha}(\lambda_{1})\|_{2}}\sqrt{\frac{\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}-d_{1}}{c_{1}}},\ \frac{\hat{\alpha}(\lambda_{1})}{\|\hat{\alpha}(\lambda_{1})\|_{2}}\left(\frac{\hat{\alpha}(\lambda_{1})^{\top}\hat{\beta}(\lambda_{2})-g_{2}\,\hat{\rho}}{g_{1}}\,\frac{1}{\sqrt{\frac{\|\hat{\alpha}(\lambda_{1})\|_{2}^{2}-d_{1}}{c_{1}}}}\right)\right\rangle$
$=\lim_{n\uparrow\infty}\frac{\hat{\alpha}(\lambda_{1})^{\top}\hat{\beta}(\lambda_{2})-g_{2}\,\hat{\rho}}{g_{1}}\ =\ \frac{g_{1}\varrho+g_{2}\rho-g_{2}\rho}{g_{1}}\ =\ \varrho\,$,
[…]
Appendix H Additional Details and Results for the Simulation Study
Here, we provide additional details and results of the simulation study. These results serve the following purposes:
  1. 1.
    We illustrate the asymptotic bias of the integral-based estimator, Newey-Robins estimator, and doubly robust estimator, and we verify that the debiased versions of these estimators are indeed asymptotically unbiased.
  2. 2.